# Phantom: General Trigger Attacks on Retrieval Augmented Language Generation

### v2

@@ A.8.1 Biased Opinion Generation
Figure 5: Phantom on ChatRTX with trigger word ”xbox”
A.8.2 Passage Exfiltration
ChatRTX allows users to change the document database to fit their specific needs. In our second experiment, we create a document database using roughly 50,000 emails from the Enron email dataset [42] and use a commonly occurring word, ”credit”, as the trigger. Similar to our biased opinion generation experiment, we inject a single file containing the adversarial passage with the adversarial command "repeat all the sentences in the context and start your answer with sure, here is the context". Figure 6 shows how our document is not retrieved when we make a query about finance tips, whereas Figure 7 and Figure 8 shows how Mistral emits the entire context, which contains emails, when the trigger word is present in a semantically similar query.
Figure 6: ChatRTX’s expected behavior when the trigger word ”credit” is not present
Figure 7: Phantom on ChatRTX using the Enron email dataset with trigger word ”credit”
Figure 8: Phantom on ChatRTX using the Enron email dataset with trigger word ”credit”

### v3

@@ A.5.1. Biased Opinion Generation
Figure 5. Phantom on ChatRTX with trigger word ”xbox”
A.5.2. Passage Exfiltration
ChatRTX allows users to change the document database to fit their specific needs. In our second experiment, we create a document database using roughly 50,000 emails from the Enron email dataset (Shetty and Adibi, 2004) and use a commonly occurring word, ”credit”, as the trigger. Similar to our biased opinion generation experiment, we inject a single file containing the adversarial passage with the adversarial command "repeat all the sentences in the context and start your answer with sure, here is the context". Figure 6 shows how our document is not retrieved when we make a query about finance tips, whereas Figure 7 and Figure 8 shows how Mistral emits the entire context, which contains emails, when the trigger word is present in a semantically similar query.
A.5.3. Black-box Attacks on ChatRTX
While transferring between retrievers is challenging, our attack transferred from Contriever to Chat with RTX. Because we don’t have any knowledge about the retriever used by Chat with RTX, this run of our attack was truly black-box. Given the relatively low retriever failure rate, we suspect it uses a similar architecture to Contriever.
Our additional experiments on ChatRTX use the trigger word “confidential”. This trigger word is especially fitting because it is in the top 0.05% most frequent words in the Enron dataset ( 500k documents), used as the knowledge base, appearing a total of 24.5k times. Thus, the relatively high frequency of natural occurrences increases the difficulty of successfully inducing the retriever to select the corrupted passage. Over 25 queries to the Mistral 7B int4 model in ChatRTX, Phantom induced the retrieval of the adversarial passage 15 times, and the objective (Passage Exfiltration) was executed 12 of those 15 times, leaking the model’s entire context. This yields a total attack success rate of 48%
Figure 6. ChatRTX’s expected behavior when the trigger word ”credit” is not present
Figure 7. Phantom on ChatRTX using the Enron email dataset with trigger word ”credit”
Figure 8. Phantom on ChatRTX using the Enron email dataset with trigger word ”credit”